FIND THE SPECIFIC VOLUME ENTHALPY INTERNAL ENERGY AND ENTROPY OF WET STEAM AT 20 BAR PRESSURE ANO DRYNESS FRACTION OF 0.8

NOTE : AT 20 BAR Values To be TAKEN FROM STEAM TABLE
1
Expert's answer
2020-06-29T08:28:04-0400

The Dryness fraction(x) of the wet steam=0.8

The pressure of wet steam(P)=20 bar=2 MPa

The saturation properties of wet steam at 2 MPa is given in the steam table as-

vf= 0.0012 m3/kg vfg= (0.0996-0.0012)=0.0984 m3/kg

hf=908.5 kJ/kg hfg= 1889.8kJ/kg

uf= 906.15 kJ/kg ufg= (2599.1-906.15) =1692.95 kJ/Kg

sf= 2.4468 kJ/kg-K sfg= 3.8923 KJ/kg-K

Specific volume at x=0.8 => v=vf+x.vfg=0.07992 m3/kg

Specific enthalpy at x=0.8 => h=hf+x.hfg= 2420.34 kJ/kg

Specific internal energy at x=0.8 =>u=uf+x.ufg=2260.51 kJ/Kg

Specific entropy at x=0.8 => s=sf+x.sfg=5.5606 kJ/kg-K